you have $121 in a savings account and save $10 per week. Your friend has $85 in a saving account and saves $16 per week. How many weeks will it take for you and you friend to have the same balance?
Answer by TimothyLamb(4379)

y = 121 + 10w
f = 85 + 16w
---
121 + 10w = 85 + 16w
6w = 36
Answer:
w = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php